# Any-Resolution AI-Generated Image Detection by Spectral Learning

### v1

@@ Abstract
Recent works have established that AI models introduce spectral artifacts into generated images and propose approaches for learning to capture them using labeled data. However, the significant differences in such artifacts among different generative models hinder these approaches from generalizing to generators not seen during training. In this work, we build upon the key idea that the spectral distribution of real images constitutes both an invariant and highly discriminative pattern for AI-generated image detection. To model this under a self-supervised setup, we employ masked spectral learning using the pretext task of frequency reconstruction. Since generated images constitute out-of-distribution samples for this model, we propose spectral reconstruction similarity to capture this divergence. Moreover, we introduce spectral context attention, which enables our approach to efficiently capture subtle spectral inconsistencies in images of any resolution. Our spectral AI-generated image detection approach (SPAI) achieves a 5.5% absolute improvement in AUC over the previous state-of-the-art across 13 recent generative approaches, while exhibiting robustness against common online perturbations.
@@ 1 Introduction
To learn a spectral model of the real images we show that the pretext task of frequency reconstruction is an effective approach for modeling their frequency distribution, under a self-supervised training setup that uses only real images. Then, to detect AI-generated images as out-of-distribution samples of this model we exploit the observation that a frequency reconstruction model trained on real images will reconstruct their frequencies more accurately, compared to the generated ones. We introduce the concept of spectral reconstruction similarity for estimating the divergence of the reconstructed frequencies from the ones actually present in the image under question.
@@
  • We introduce the spectral distribution of real images as a suitable invariant pattern for distinguishing between real and AI-generated images and we propose masked spectral learning for modeling it.
@@ 2 Related Work
Figure 2: Overview of the SPAI approach. We learn a model of the spectral distribution of real images under a self-supervised setup using masked spectral learning. Then, we use the spectral reconstruction similarity to measure the diverge from the learned distribution and detect synthetic images as out-of-distribution samples of this model, while spectral context vector captures the spectral context under which the spectral reconstruction similarity values are computed. Finally, spectral patch attention enables the processing of any-resolution images for capturing subtle spectral differences.
AI-Generated Image Detection: In response, several approaches have been proposed for distinguishing real from AI-generated (synthetic) images, either by detecting semantic inconsistencies or by capturing low-level artifacts introduced by the generative models. In the former group, recent approaches detect inconsistencies in facial geometry [3] as well as in shadows [61], perspective [21] and lighting [20] of the image. However, as improved generative models get released, such visible inconsistencies cease to exist. To this end, recent works have established that generative models introduce low-level artifacts and proposed approaches for capturing them either in the spatial [71, 7] or in the spectral [23, 1, 45] domain. However, such artifacts significantly differ even between models with minimal differences [1]. In pursuit of artifacts that better generalize across different generators, others have proposed architectures for capturing inconsistencies in texture-rich image regions [48, 34, 77], gradient differences [65] or artifacts introduced by the upscaling layers of the generative models [66] as well as spectral augmentation approaches [15]. Moreover, several recent works have employed features from pre-trained CLIP [56] encoders for capturing both low-level and semantic artifacts [39, 8, 52, 63]. Common ground in all the aforementioned works is the attempt to model artifacts introduced by the generative approaches, with the intention of using them for the AID task. As such, they fail to generalize on unseen generators that introduce artifacts different than the ones present in the training set.
@@ 3.1 Masked Spectral Learning
Frequency Masking: To obtain the low- and high-frequency components of input image $x\in\mathbb{R}^{H\times W}$, we use the 2D Discrete Fourier Transform $\chi=\mathcal{F}(x)$, where $\chi(u,v)$ defines the complex frequency value at coordinates $(u,v)$. Next, inspired by [73] we define a mask $M$ for masking parts of the spectrum according to a fixed radius $r$ from its center $(c_{H},c_{W})$, with $d(\cdot,\cdot)$ denoting the Euclidean distance:
@@
To train $\mathcal{G}$ we use an auxiliary decoding head $\mathcal{H}$ to predict from its output the reconstructed image $\hat{x}=\mathcal{H}(\mathcal{G}(\mathcal{B}(x^{l},x^{h})))$, where $\mathcal{B}(\cdot,\cdot)$ randomly samples one of its input according to a binomial distribution. We train the combination of $\mathcal{G}$ and $\mathcal{H}$ using the objective $\mathcal{L}_{rec}=\mathcal{D}(\mathcal{F}(x),\mathcal{F}(\hat{x}))$, where $\mathcal{D}$ defines the frequency distance introduced in [73]. $\mathcal{H}$ is discarded after training.
@@ 3.2 Spectral Reconstruction Similarity
Up to this point, $\mathcal{G}$ has learned to reconstruct the spectral distribution of real images given their low- or high-frequency components. Here we introduce an approach to detect the out-of-distribution samples with regard to $\mathcal{G}$. In the rest of the text, we assume that $\mathcal{G}$ is a Vision Transformer (ViT) [16].
@@
We encode $x$, $x^{h}$ and $x^{l}$ into $z_{n}$, $z_{n}^{h}$ and $z_{n}^{l}$ respectively, using Eq. 4. Then, we use $N$ projection operators, denoted $\mathbf{P}_{n}:R^{d}\rightarrow R^{D}$, one per block of $\mathcal{G}$, to project all three aforementioned groups of representations from the feature space of $\mathcal{G}$ to one that facilitates the similarity operations that we define next. Each of these projection operators is composed of a series of linear projection, layer normalization [43] and GELU activation layers [31]. Further details regarding their structure are provided in the supplementary material. We use them to project the representations of $z_{n}$, $z_{n}^{h}$ and $z_{n}^{l}$ into $\bar{z}_{n}$, $\bar{z}_{n}^{h}$ and $\bar{z}_{n}^{l}$, all belonging to the space of $\mathbb{R}^{L\times D}$. In the rest of our approach we only use these latter representations.
As $\mathcal{G}$ constitutes a spectral model of real images, it is expected to better reconstruct their missing frequencies compared to the ones of AI-generated images. Therefore, we expect bigger distances between $\bar{z}_{n}$, $\bar{z}_{n}^{h}$ and $\bar{z}_{n}^{l}$ for AI-generated images. As such we expect the similarity among the features of the original image, the low- and the high-frequency components to indicate the alignment of the spectral distribution of an image with the learned distribution of $\mathcal{G}$. To measure this divergence we introduce the concept of Spectral Reconstruction Similarity (SRS). For two representations $z^{A}$ and $z^{B}$ in space $\mathbb{R}^{L\times D}$ we define SRS as:
@@ 3.3 Spectral Context Vector
Next, we define a learnable spectral map $C\in\mathbb{R}^{N\times D}$ and two projection functions $\mathcal{P}_{1}(\cdot):R^{2D}\rightarrow R^{D}$ and $\mathcal{P}_{2}(\cdot):R^{D}\rightarrow R^{D}$. Further details regarding the structure of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are provided in supplementary material. Using these three components the network learns to construct the SCV by attending to the most useful of the $D$ feature values for each of the $N$ transformer blocks of $\mathcal{G}$ as following:
@@ 3.4 Spectral Context Attention
In practice, the height and width of the input image $x$ may receive arbitrarily large values, i.e. images of many megapixels. While ViTs in theory can scale to inputs of arbitrary size, their quadratic computational complexity w.r.t the length of their token sequence practically prohibits their use on such large inputs. Similar limitations are faced by convolutional neural networks. Thus, a typical approach in most computer vision tasks that rely on the semantic information of the image is to resize it to a fixed size. However, in verification tasks, such as AID, the ability to capture anomalies in subtle details of the image is crucial [7]. So, a resizing operation, that effectively discards the high-frequency information of the image, is problematic. To tackle this issue we introduce the concept of Spectral Context Attention (SCA), for combining the most discriminative SRS values computed for different patches of the image, according to their respective SCV and a learnable spectral reconstruction importance vector $q\in\mathbb{R}^{D_{h}}$.
Initially, we split the original image into $K$ patches of size $h\times w$, forming the patched representation of the image $x^{P}\in\mathbb{R}^{K\times h\times w}$, while $x_{k}$ denotes the $k$-th patch of this sequence. For each patch, we construct its spectral vector $z_{k}^{S}$ following the approach introduced throughout the previous paragraphs.
Next, following the transformer’s attention notation [70] we define the weight matrices $W_{K},W_{V}\in\mathbb{R}^{(D+6N)\times D_{h}}$ and $W_{O}\in\mathbb{R}^{D_{h}\times(D+6N)}$. Conveniently, we define $\bar{z}^{S}\in\mathbb{R}^{K\times(D+6N)}$ as the concatenation of spectral vectors for all $K$ image patches. Then, we fuse them into the image-level spectral vector $\mathbf{z}^{S}$ through the following:
@@ 3.5 Training Process and Implementation
Spectral Model: As our spectral model of real images $\mathcal{G}$ we use a ViT-B/16 transformer [16] pre-trained on ImageNet by Xie et al. [73], using a masking radius $r=16$. We keep its weights frozen throughout the training of the components introduced in Secs. 3.2, 3.3 and 3.4.
@@ 4.1 Evaluation Setup
Implementation Details: We train our architecture for 35 epochs, with a learning rate ($lr$) of $45e-$ 4, cosine decay and five epochs of linear warmup, using the AdamW [49] optimizer and batch size of 72. We set $D=1024$ and, to align with the pretext task, $r=16$. The number of transformer blocks in the ViT we use for $\mathcal{G}$ is $N=12$, its latent dimensionality is $d=768$ while the size of each of the $K$ patches is $h=w=224$ pixels and empirically we set $D_{h}=1536$. We train our model by applying random resizing, cropping, rotation, Gaussian blur, Gaussian noise and JPEG compression augmentations. The training of the SRS, SCV and SCA modules is performed on a single Nvidia L40S 48GB GPU.
@@
Datasets: Following [39, 7], we train our approach using the training data by Corvi et al. [7] that include 180k latent diffusion [59] images as well as 180k real images, all of size $256\times 256$. Then, to evaluate our approach, we compose a representative set of images originating from recently introduced generative models. In particular, we employ images from [1] for Glide [51], Stable Diffusion 1.3 (SD1.3) and Stable Diffusion 1.4 (SD1.4), representing some early popular text-to-image synthesis approaches, primarily generating low-resolution images, i.e. $<0.5$ megapixels. Next, we collect from [1] images for DALLE2 [57], Stable Diffusion 2 (SD2) and Stable Diffusion XL (SDXL) [55]. Also, we employ the online sources of [42, 24] to collect images generated by Flux and Stable Diffusion 3 (SD3) [19] respectively. We also include images generated and provided by GigaGAN [35], which constitutes one of the few recently proposed GAN-based generators. These approaches represent a group of recent methods, capable of producing images up to 1 megapixel. Finally, we collect images for Midjourney-v5 (MJv5), DALLE3 [2] and Firefly from [1], as well as for Midjourney-v6.1 (MJv6.1) from [60]. For each model we collect 1k images, with the exception of MJv6.1 for which we use 631 publicly available images. Regarding the real images we employ 1k samples from each of RAISE [11] and FODB [29], as representative sources of images captured by DSLR cameras and smartphones respectively. We include another 1k images from the test sets of ImageNet [14], COCO [47] and Open Images [41] that represent a general collection of online images, an object-focused database and a collection of high-resolution Web images respectively. These datasets have been collected before the advent of capable generative AI approaches and thus can be safely assumed to be real.
Figure 3: Robustness evaluation on common perturbations. Average AUC is presented over the perturbed versions of two sources of authentic images from smartphones and DSLR cameras respectively and 13 generative models.
@@ 4.2 Comparison Against State-of-the-Art
We compared our approach against all the considered AID methods and we report the obtained AUC scores for each of the 13 considered generative models in Tab. 1, averaged over all the sources of real images. While almost all competitive methods successfully distinguish AI-generated images originating from specific generators, in some cases even achieving near-perfect performance, they catastrophically fail to others. For example, while in the case of Firefly images our method achieves the best performance, in the case of DALLE3 images NPR [66] performs best. However, NPR underperforms in images from most of the other generative approaches, thus scoring the worst on average among the considered detectors. This highlights the inability of AID approaches that learn to capture specific types of artifacts to generalize to generators that introduce different types of inconsistencies. Instead, our approach, exhibits consistently high detection performance across all the generative models, achieving an absolute average improvement of 5.5% over the second best method.
Robustness Against Online Perturbations: To understand the robustness of our network against common perturbations encountered online we study five different types of perturbations. In particular, we study the effects of JPEG and WebP compression with quality factors $Q={85,70,50}$ as well as Gaussian Blur with kernel size $k={3,5,7}$, Gaussian Noise with standard deviation $\sigma={1,3,5}$ and resizing with scaling factors of ${85,70,50}\%$. To better isolate the effect of each perturbation, in this study we use only the sets of real images originating directly from DSLR and smartphone cameras, i.e. from RAISE and FODB respectively, and present the results in Fig. 3. For illustration purposes we limit the results to the two best competitive approaches, as indicated by Tab. 1. We see that in all cases our approach outperforms the competing methods.

### v2

@@ Abstract
Recent works have established that AI models introduce spectral artifacts into generated images and propose approaches for learning to capture them using labeled data. However, the significant differences in such artifacts among different generative models hinder these approaches from generalizing to generators not seen during training. In this work, we build upon the key idea that the spectral distribution of real images constitutes both an invariant and highly discriminative pattern for AI-generated image detection. To model this under a self-supervised setup, we employ masked spectral learning using the pretext task of frequency reconstruction. Since generated images constitute out-of-distribution samples for this model, we propose spectral reconstruction similarity to capture this divergence. Moreover, we introduce spectral context attention, which enables our approach to efficiently capture subtle spectral inconsistencies in images of any resolution. Our spectral AI-generated image detection approach (SPAI) achieves a 5.5% absolute improvement in AUC over the previous state-of-the-art across 13 recent generative approaches, while exhibiting robustness against common online perturbations. Code is available on https://mever-team.github.io/spai.
@@ 1 Introduction
To learn a spectral model of the real images we show that the pretext task of frequency reconstruction is an effective approach for modeling their frequency distribution, under a self-supervised training setup that uses only real images. Then, to detect AI-generated images as out-of-distribution samples of this model, we exploit the observation that a frequency reconstruction model trained on real images will reconstruct their frequencies more accurately, compared to the generated ones. We introduce the concept of spectral reconstruction similarity for estimating the divergence of the reconstructed frequencies from the ones actually present in the image under question.
@@
  • We introduce the spectral distribution of real images as a suitable invariant pattern for distinguishing between real and AI-generated images and we propose the use of masked spectral learning for modeling it.
@@ 2 Related Work
Figure 2: Overview of the SPAI approach. We learn a model of the spectral distribution of real images under a self-supervised setup using masked spectral learning. Then, we use the spectral reconstruction similarity to measure the divergence from this learned distribution and detect AI-generated images as out-of-distribution samples of this model. Spectral context vector captures the spectral context under which the spectral reconstruction similarity values are computed, while spectral context attention enables the processing of any-resolution images for capturing subtle spectral inconsistencies.
AI-Generated Image Detection: In response, several approaches have been proposed for distinguishing real from AI-generated (synthetic) images, either by detecting semantic inconsistencies or by capturing low-level artifacts introduced by the generative models. In the former group, recent approaches detect inconsistencies in facial geometry [3] as well as in shadows [62], perspective [22] and lighting [21] of the image. However, as improved generative models get released, such visible inconsistencies cease to exist. To this end, recent works have established that generative models introduce low-level artifacts and proposed approaches for capturing them either in the spatial [72, 7] or in the spectral [18, 1, 46] domain. However, such artifacts significantly differ even between models with minimal differences [1]. In pursuit of artifacts that better generalize across different generators, others have proposed architectures for capturing inconsistencies in texture-rich image regions [49, 35, 79], gradient differences [66] or artifacts introduced by the upscaling layers of the generative models [67] as well as spectral augmentation approaches [15]. Moreover, several recent works have employed features from pre-trained CLIP [57] encoders for capturing both low-level and semantic artifacts [40, 8, 53, 64, 76]. Common ground in all the aforementioned works is the attempt to model artifacts introduced by the generative approaches, with the intention of using them for the AID task. As such, they fail to generalize on unseen generators that introduce artifacts different than the ones present in the training set.
@@ 3.1 Masked Spectral Learning
Frequency Masking: To obtain the low- and high-frequency components of input image $x\in\mathbb{R}^{H\times W}$, we use the 2D Discrete Fourier Transform $\chi=\mathcal{F}(x)\in\mathbb{R}^{H\times W}$, where $\chi(u,v)$ defines the complex frequency value at coordinates $(u,v)$. Next, inspired by [74], we define a mask $M$ for masking parts of the spectrum according to a fixed radius $r$ from its center, that we denote as $(c_{H},c_{W})$, with $d(\cdot,\cdot)$ denoting the Euclidean distance:
@@
To train $\mathcal{G}$ we use an auxiliary decoding head $\mathcal{H}$ to predict from its output the reconstructed image $\hat{x}=\mathcal{H}(\mathcal{G}(\mathcal{B}(x^{l},x^{h})))$, where $\mathcal{B}(\cdot,\cdot)$ randomly samples one of its inputs according to a binomial distribution. We train the combination of $\mathcal{G}$ and $\mathcal{H}$ using the objective $\mathcal{L}_{rec}=\mathcal{D}(\mathcal{F}(x),\mathcal{F}(\hat{x}))$, where $\mathcal{D}$ defines the frequency distance introduced in [74]. $\mathcal{H}$ is discarded after training.
@@ 3.2 Spectral Reconstruction Similarity
Up to this point, $\mathcal{G}$ has learned to reconstruct the spectral distribution of real images given their low- or high-frequency components. Here, we introduce an approach to detect the out-of-distribution samples of $\mathcal{G}$. In the rest of the text, we assume that $\mathcal{G}$ is a Vision Transformer (ViT) [16].
@@
We encode $x$, $x^{h}$ and $x^{l}$ into $z_{n}$, $z_{n}^{h}$ and $z_{n}^{l}$ respectively, using Eq. 4. Then, we use $N$ projection operators, denoted $\mathbf{P}_{n}:R^{d}\rightarrow R^{D}$, one per block of $\mathcal{G}$, to project the three aforementioned groups of representations from the feature space of $\mathcal{G}$ to one that facilitates the similarity operations that we define next. Each projection operator comprises a series of linear projection, layer normalization [44] and GELU activation layers [32]. Further details regarding their structure are provided in the supplementary material. We use them to project $z_{n}$, $z_{n}^{h}$ and $z_{n}^{l}$ into $\bar{z}_{n}$, $\bar{z}_{n}^{h}$ and $\bar{z}_{n}^{l}$, all belonging to the space of $\mathbb{R}^{L\times D}$. In the rest of our approach we only use these latter representations.
As $\mathcal{G}$ constitutes a spectral model of real images, it is expected to better reconstruct their missing frequencies compared to the ones of AI-generated images. Therefore, we expect bigger distances between $\bar{z}_{n}$, $\bar{z}_{n}^{h}$ and $\bar{z}_{n}^{l}$ for AI-generated images. As such, we expect the similarity among the features of the original image, the low- and the high-frequency components to indicate the alignment of the spectral distribution of an image with the learned distribution of $\mathcal{G}$. To measure this divergence we introduce the concept of Spectral Reconstruction Similarity (SRS). For two representations $z^{A}$ and $z^{B}$ in space $\mathbb{R}^{L\times D}$ we define SRS as:
@@ 3.3 Spectral Context Vector
Next, we define a learnable spectral map $C\in\mathbb{R}^{N\times D}$ and two projection functions $\mathcal{P}_{1}(\cdot):R^{2D}\rightarrow R^{D}$ and $\mathcal{P}_{2}(\cdot):R^{D}\rightarrow R^{D}$. Further details regarding the structure of $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ are provided in supplementary material. Using these three components the network learns to construct the SCV by attending to the most useful of the $2D$ feature values for each of the $N$ transformer blocks of $\mathcal{G}$ as following:
@@ 3.4 Spectral Context Attention
In practice, the height and width of the input image $x$ may receive arbitrarily large values, i.e. images of many megapixels. While ViTs in theory can scale to inputs of arbitrary size, their quadratic computational complexity w.r.t the length of their token sequence practically prohibits their use on such large inputs. Similar limitations are faced by convolutional neural networks. Thus, a typical approach in most computer vision tasks that rely on the semantic information of the image is to resize it to a fixed size. However, in verification tasks, such as AID, the ability to capture anomalies in subtle details of the image is crucial [7]. So, a resizing operation, that effectively discards the high-frequency information of the image, is problematic. To tackle this issue, we introduce the concept of Spectral Context Attention (SCA), for combining the most discriminative SRS values computed for different patches of the image, according to their respective SCV and a learnable spectral reconstruction importance vector $q\in\mathbb{R}^{D_{h}}$.
Initially, we split the original image into $K$ patches of size $h\times w$, forming the patched representation of the image $x^{P}\in\mathbb{R}^{K\times h\times w}$, while $x_{k}$ denotes the $k$-th patch of this sequence. For each patch, we construct its spectral vector $z_{k}^{S}$ following the approach introduced throughout the previous paragraphs. Next, following the transformer’s attention notation [71] we define the weight matrices $W_{K},W_{V}\in\mathbb{R}^{(D+6N)\times D_{h}}$ and $W_{O}\in\mathbb{R}^{D_{h}\times(D+6N)}$. Conveniently, we define $\bar{z}^{S}\in\mathbb{R}^{K\times(D+6N)}$ as the concatenation of spectral vectors for all $K$ image patches. Then, we fuse them into the image-level spectral vector $\mathbf{z}^{S}$ through the following:
@@ 3.5 Training Process and Implementation
Spectral Model: As our spectral model of real images $\mathcal{G}$ we use a ViT-B/16 transformer [16] pre-trained on ImageNet [14] by Xie et al. [74], using a masking radius $r=16$. We keep its weights frozen throughout the training of the components introduced in Secs. 3.2, 3.3 and 3.4.
@@ 4.1 Evaluation Setup
Implementation Details: We train our architecture for 35 epochs, with a learning rate ($lr$) of $45e-$ 4, cosine decay and five epochs of linear warmup, using the AdamW [50] optimizer and batch size of 72. We set $D=1024$, $K_{training}=4$ and, to align with the pretext task, $r=16$. The number of transformer blocks in the ViT we use for $\mathcal{G}$ is $N=12$, its latent dimensionality is $d=768$, while the size of each of the $K$ patches is $h=w=224$ pixels and empirically we set $D_{h}=1536$. We train our model by applying random resizing, cropping, rotation, Gaussian blur, Gaussian noise and JPEG compression augmentations. The SRS, SCV and SCA are trained on a single Nvidia L40S 48GB GPU.
@@
Datasets: Following [40, 7], we train our approach on the 180k low resolution images of size $256\times 256$, from a single latent diffusion model [60], introduced along with 180k real images by Corvi et al. [7]. Then, to evaluate our approach, we compose a representative set of images originating from recently introduced generative models. In particular, we employ images from [1] for Glide [52], Stable Diffusion 1.3 (SD1.3) and Stable Diffusion 1.4 (SD1.4), representing some early popular text-to-image synthesis approaches, primarily generating low-resolution images, i.e. $<0.5$ megapixels. Next, we collect from [1] images for DALLE2 [58], Stable Diffusion 2 (SD2) and Stable Diffusion XL (SDXL) [56]. Also, we employ the online sources of [43, 25] to collect images generated by Flux and Stable Diffusion 3 (SD3) [20] respectively. We also include images generated and provided by GigaGAN [36], which constitutes one of the few recently proposed GAN-based generators. These approaches represent a group of recent methods, capable of producing images up to 1 megapixel. Finally, we collect images for Midjourney-v5 (MJv5), DALLE3 [2] and Firefly from [1], as well as for Midjourney-v6.1 (MJv6.1) from [61]. For each model we collect 1k images, with the exception of MJv6.1 for which we use 631 publicly available images. Regarding the real images we employ 1k samples from each of RAISE [11] and FODB [30], as representative sources of images captured by DSLR cameras and smartphones respectively. We include another 1k images from the test sets of ImageNet [14], COCO [48] and Open Images [42] that represent a general collection of online images, an object-focused database and a collection of high-resolution Web images respectively. These datasets have been collected before the advent of capable generative AI approaches and thus can be safely assumed to be real.
@@
Figure 3: Robustness evaluation on common perturbations. Average AUC is presented over the perturbed versions of two sources of real images from smartphones and DSLR cameras respectively and 13 generative models.
@@ 4.2 Comparison Against State-of-the-Art
We compare against the considered detectors and report the obtained AUC scores for each of the 13 considered generative models in Tab. 1, averaged over all the sources of real images. While almost all competitors successfully distinguish AI-generated images from specific generators, in some cases even achieving near-perfect performance, they catastrophically fail to others. For example, while in the case of Firefly images our method achieves the best performance, in the case of DALLE3 images NPR [67] performs best. However, NPR underperforms in all the rest generative approaches, overall scoring the worst among the considered detectors. This highlights the inability of detectors that learn to capture specific types of artifacts to generalize to generators that introduce different types of inconsistencies. Instead, our approach, exhibits consistently high detection performance across all the generative models, achieving an absolute average improvement of 5.5% over the second best method, even with the later using a much larger ViT-L backbone pre-trained on the 400 million images of CLIP.
Robustness Against Online Perturbations: To understand the robustness of our network against common perturbations encountered online we study five different types of perturbations. In particular, we study the effects of JPEG and WebP compression with quality factors $Q={85,70,50}$ as well as Gaussian Blur with kernel size $k={3,5,7}$, Gaussian Noise with standard deviation $\sigma={1,3,5}$ and resizing with scaling factors of ${85,70,50}\%$. To better isolate the effect of each perturbation, in this study we use only the sets of real images originating directly from DSLR and smartphone cameras, i.e. from RAISE and FODB respectively, and present the results in Fig. 3. For illustration purposes we limit the results to the two best competitor approaches, as indicated by Tab. 1. We see that in all cases our approach outperforms the competing methods.